OpenStudy (anonymous):

Stan can mow a lawn in 3 hrs while Scott can do the same in 2 hrs. Hw long it would take them together to mow the lawn?

OpenStudy (anonymous):

6/5 = 1.2 hrs

OpenStudy (anonymous):

Call y=the total work done and x=total work done per hour. Stan does: y=3x and Scott does: y=2x Together they do the sum of the two equations, 2y=5x or, y=(5/2)x So they can get the work done in 2.5 hours.

OpenStudy (anonymous):

Wait, this is wrong. If two Stan's were mowing the lawn they would finish in 1.5 hours. My calculation would yield 3hours in this case. We need to divide (5/2) by 2. The answer should be (5/4)=1.25 hours

OpenStudy (anonymous):

eseidl.. I want to contradict your answer.. When Scott can complete within 2 hrs, so if Stan joins him then doesn't the job should be complete within 2 hrs?

OpenStudy (anonymous):

If Stan can complete in 3 hrs, then job done by him in 1 hr is 1/3 If Scott can complete in 2 hrs, then job done by him in 1 hr is 1/2 Together, they complete the work of 1/3 + 1/2 in 1 hr 1/3 + 1/2 = 5/6 So, together they complete 5/6th of work in 1 hr. Hence, work is completed in 6/5 hrs = 1.2 hrs...

OpenStudy (anonymous):

My line of reasoning goes as follows. Stan and scott can complete the task in 3hours and 2 hours respectively. On average they can complete the task in 2.5 hours. If they work on it together, they both work at this average rate of 2.5 hours but since there is two individually doing the work it takes half the time, i.e., 1.25 hours.

OpenStudy (anonymous):

Eseidl.. I am sorry to say this.. but let us have a check by the third person...

OpenStudy (anonymous):

Actually I think you may be correct.

OpenStudy (anonymous):

I "MAY" be correct? Let's have an expert opinion.. If I am wrong then I accept and correct my mistake..

OpenStudy (anonymous):

hmmmm yeah I'd like a third opinion.

OpenStudy (anonymous):

i think helpingtutors is correct

OpenStudy (anonymous):

The problem with Tutor's reasoning is that it relies on Stan and Scott's work rates while they work in isolation. This is NOT the same in general as their work rates when they work together. Scott will be slowed down slightly when he works with Stan. Averaging their rates of work corrects for this.

OpenStudy (anonymous):

eseidl... we have 3rd person's opinion(mathematicalpro). would you like one more opinion?

OpenStudy (anonymous):

He didn't seem 100% sure, and he provided no argument with his opinion.

OpenStudy (anonymous):

Yes you are right.. Let us wait for another opinion,,,

OpenStudy (anonymous):

agreed. btw, are you sure you are correct (I'm not sure I am, by the way lol)?

OpenStudy (akshay_budhkar):

I am 100 percent sure that the answer is 6/5=1.2... Now i don't know which of you is correct :)

OpenStudy (akshay_budhkar):

I believe helping tutor is correct

OpenStudy (anonymous):

x/2+x/3=1 2x+3x/6=1 5x=6 x=6/5 x=1.2

OpenStudy (anonymous):

1.2 hours it is. thanks guys!